joe is 3/5 as old as brother sam. in 5 years sam will be one and one half times as old as joe.. how old is joe now.
Answer by mananth(16946)

Sam age ==x
Joe age = (3/5)x

after 5 years
sam = x+5 years
Joe = ((3/5)x+5)
x+5 = (3/2)*(3/5x+5)
multiply by 2
2(x+5)=3((3/5)x+5)
2x+10 = 3(3x+25)/5
multiply by 5
10x+50=9x+75
10x-9x=75-50
x=25 sam's age
Joes age 3/5 * 25=15 years
m.ananth@hotmail.ca